a sample of hydrogen at 37.0 C exerts pressure of 422 torrr. If the gas is heated to 63.) C at constant volume, what is the new pressure of hydrogen?


1
Expert's answer
2021-03-03T01:58:48-0500

P/T = Const, P1/T1 = P2/T2, P2 = P1*T2/T1 = 422*(63+273.15)/(37+273.15) = 457.4 torr.